OpenStudy (anonymous):

PLEASE HELP: 3x - 2y = 5 -6x + 4y = 7

OpenStudy (anonymous):

Oh, solve the system with ELIMINATION!!

OpenStudy (jdoe0001):

so, what do you think from the previous explanation by mathstudent55 then?

OpenStudy (akashdeepdeb):

In the first equation 3x = 5 + 2y or x = (5+2y)/3 Now put this x value in the second equation! :D

OpenStudy (anonymous):

@jdoe0001 I think it would be no solution.. because it would be 0=7. Idk if I'm right.

OpenStudy (jdoe0001):

you're correct, because both cancel out

OpenStudy (jdoe0001):

3x - 2y = 5 -6x + 4y = 7 --------------------- \(\bf y = \cfrac{3}{2}x -5\\ y = \cfrac{3}{2}x+7\) notice the slopes, thus they're parallel, they never touch, thus no solution

OpenStudy (jdoe0001):

well, \(\bf y = \cfrac{3}{2}x -5\\ y = \cfrac{3}{2}x+\cfrac{7}{4}\) missed something before, but the slopes are the same, and thus it's so

OpenStudy (jdoe0001):

ack, more typos

OpenStudy (jdoe0001):

\(\bf y = \cfrac{3}{2}x -\cfrac{5}{2}\\ y = \cfrac{3}{2}x+\cfrac{7}{4}\)

OpenStudy (anonymous):

@jdoe0001 omg, i see it now, thank you so much. everyone elses replies confused me. thank you math god/goddess

OpenStudy (jdoe0001):

yw

OpenStudy (anonymous):

i might have alot more questions for you. how can i make sure you answer them?

OpenStudy (anonymous):

btw, just correcting my self it would be 0+0 = 12, which is not a solution. RIGHTT?